OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (2,6) and (-4, 8). Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (aripotta):

what's the first thing you have to do?

OpenStudy (anonymous):

Honestly clueless

OpenStudy (aripotta):

you have to first find the slope

OpenStudy (anonymous):

I just need this answered to be done with my online ALG 2 class

OpenStudy (aripotta):

i'm in algebra 2 as well. just started though

OpenStudy (anonymous):

sooo slope = -1/3?

OpenStudy (aripotta):

yea, so now what do we do?

OpenStudy (anonymous):

sub in coordinate points in the equation i guess

OpenStudy (aripotta):

yea, we have to put the information now into point-slope form. do you know what that is?

OpenStudy (anonymous):

Y=MX+B right

OpenStudy (aripotta):

no, that's slope-intercept form

OpenStudy (aripotta):

from point-slope form, we can change it into slope-intercept, and then standard

OpenStudy (anonymous):

erp. whats point slope?

OpenStudy (aripotta):

y - y1 = m(x - x1)

OpenStudy (aripotta):

so we can use one of the given coordinates for this. we'll use the first one (2,6)

OpenStudy (aripotta):

you substitute the 2 for x1, 6 for y1, and -1/3 for m

OpenStudy (aripotta):

is this making sense so far?

OpenStudy (anonymous):

x-2 = -1/3(y-6) ?

OpenStudy (aripotta):

you switched up the x's and y's point-slope is: y - y1 = m(x - x1) so it would be y - 6 = -1/3(x - 2)

OpenStudy (anonymous):

typing error :/ okay now what

OpenStudy (aripotta):

so now it's in point-slope...how could we get it into slope-intercept form? y = mx + b?

OpenStudy (anonymous):

add 6 to both sides?

OpenStudy (aripotta):

well first, i'd distribute the -1/3 into the parentheses, then add 6

OpenStudy (aripotta):

so what would it be in slope-intercept form?

OpenStudy (anonymous):

the 1/3 is messing me up trying to distribute

OpenStudy (aripotta):



OpenStudy (anonymous):

-2/3 right?

OpenStudy (anonymous):

positive my bad

OpenStudy (aripotta):

yep

OpenStudy (aripotta):

so what's 2/3 + 6?